1a)Calculate the pressure exerted by 1 mole of methane in a 250ml container at 300k using vanderwaal's equation

1b) what pressure will be predicted by ideal gas equation

a=2.253L^2atm/mol^2

b=0.0428L/mol

R=0.0821Latm/mol/k



1
Expert's answer
2021-07-16T01:29:20-0400



Volume = "\frac{250}{1000}=0.25l"



"[P+\frac{n^2a}{V^2}][V-nb]=nRT"


"(P+\frac{2.253}{0.25×0.25})(0.25-0.0428)=0.0821×300"


"(P+36.048)(0.2072)=24.63"


"P+36.048=118.87"


"P=82.82" atm